OpenStudy (anonymous):

A rectangular box is twice as long as it is wide. The height of the box is 3 feet less than the width. If the box is x feet wide, what polynomial represents its volume in cubic feet? the answer is NOT 2x2 - 6x

OpenStudy (ganpat):

w = x l = 2x h = x-3 V = x (2x) (x -3) V = 2x2 (x -3) V = 2x3 - 6x2...